four years ago, jonny invested $4000 at 5.4% intrest compunded annually. five years ago, he invested $3700 at 5.6% intrest compounded quartly. when are the two investments worth the same (or has the smaller one alwreday surprassed the greater one)?
Answer by stanbon(75887)

four years ago, johnny invested $4000 at 5.4% interest compounded annually.
----
five years ago, he invested $3700 at 5.6% interest compounded quarterly.
====
when are the two investments worth the same (or has the smaller one already surpassed the greater one)?
------
Value of the $3700 investment after one year.
3700(1+0.056) = $3907
------
Solve for "t":
3907(1+0.056)^t = 4000(1+0.054)^t
[1.056/1.054]^t = 4000/3907
1.001898^t = 1.02375
----
t = log(1.02375)/log(1.001898)
---
t = 12.38 years
======================
Same value 8 1/3 years from now.
======================
Cheers,
Stan H.